OpenStudy (anonymous):

both logs have base 2. log(X)^2=(log(x))^2 I have never come across a question where I need to square the log can anybody help

OpenStudy (mathstudent55):

\(\log_2 x^2 = \log_2^2 x\) \(2 \log_2 x = \log_2^2 x\) \(\log_2^2 x - 2 \log_2 x = 0\) \(\log_2 x(\log_2 x - 2) =0\) \(\log_2 x = 0 \) or \(\log_2x = 2\) \(x = 1\) or \(x = 4\)

OpenStudy (mathstudent55):

Solve it as a quadratic in \(\log_2 x\). Then solve the log parts for x.